### v1

# Boundary Effects and Oxygen Deficiency-Driven Pattern Transitions in Algal Bioconvection

Abstract
Suspensions of motile microorganisms can spontaneously give rise to large-scale fluid motion, known as bioconvection, which is characterized by dense, cell-rich downwelling plumes interspersed with broad upwelling regions. In this study, we investigate bioconvection in shallow suspensions of Chlamydomonas reinhardtii cells confined within spiral-shaped boundaries, combining detailed experimental observations with three-dimensional simulations. Under open liquid–air interface conditions, cells accumulate near the surface due to negative gravitaxis, forming spiral-shaped density patterns that subsequently fragment into lattice-like structures and give rise to downwelling plumes. Space–time analyses reveal coherent rotational dynamics, with inward-moving patterns near the spiral core and bidirectional motion farther from the center. Introducing confinement by sealing the top boundary with an air-impermeable transparent wall triggers striking transitions in the bioconvection patterns, driven by oxygen depletion: initially stable structures reorganize into new patterns with reduced characteristic wavelengths. Further eliminating the fluid gap above the spiral boundary extends the duration of this transitional phase, highlighting the essential role of fluid-mediated cell-to-cell communication in coordinating collective behavior. Complementary 3D simulations, based on the incompressible Navier–Stokes equations and incorporating negative buoyancy and active stress from swimming cells, capture the initial pattern formation and its subsequent instability, reproducing the fragmentation of spiral-shaped accumulations into downwelling plumes and the emergence of strong vortical flows—nearly an order of magnitude faster than individual cell swimming speeds. However, these models do not capture the oxygen-driven pattern transitions observed experimentally. Our findings reveal that confinement geometry, oxygen dynamics, and metabolic transitions critically govern bioconvection pattern evolution, offering new strategies to control microbial self-organization and flow through environmental and geometric design.
  • These authors contributed equally to this work.
Keywords: Chlamydomonas reinhardtii, Bioconvection, Micro-swimmers, Hydrodynamic instabilities, Aerobic to anaerobic metabolism transition.
1 Introduction
Microswimmers such as Escherichia coli, Bacillus subtilis, and the freshwater alga Chlamydomonas reinhardtii generate characteristic fluid flows through the coordinated beating of their flagella and body movements [1, 2, 3, 4, 5, 6, 7]. Depending on their flagellar arrangement, these microorganisms are typically categorized as either ”pushers” or ”pullers” [8, 9, 4, 10, 11]. In pusher-type swimmers, such as E. coli, the thrust generated by flagellar motion propels fluid outward along the swimming axis and draws it inward from the sides, creating extensile flow fields. Conversely, puller-type swimmers, such as Chlamydomonas reinhardtii, use anterior flagella to pull themselves forward, drawing fluid inward along the swimming axis and pushing it outward laterally, producing contractile flow fields. These microscale hydrodynamic disturbances significantly influence the motion of nearby passive particles and neighboring microorganisms, leading to complex collective behaviors, especially in dense suspensions. In particular, when large populations of motile microorganisms interact through the flows they generate, they can spontaneously organize into coherent large-scale structures, producing macroscopic flows through a process known as bioconvection. Unlike thermally driven convection, bioconvection arises solely from the active motility and orientation biases of the cells [9, 12, 13, 14, 15, 16, 17], and is characterized by the emergence of dense, cell-rich downwelling plumes interspersed with broad regions of upwelling fluid.
Bioconvection is commonly observed in shallow laboratory suspensions, where cells swim upward on average due to directional responses known as ”taxes,” including gravitaxis (orientation against gravity), gyrotaxis (a balance between gravitational and viscous torques), phototaxis (response to light), and chemotaxis (response to chemical gradients) [18, 19, 15, 19, 9, 12, 20, 21, 22]. Gravitaxis, in particular, is typically a passive orientation mechanism resulting from physical asymmetry within the cells, such as bottom-heavy mass distribution [12, 23, 24]. Cells swimming upward due to negative gravitaxis accumulate near the upper suspension boundary, creating an unstable density gradient that triggers convective instabilities and leads to the emergence of bioconvection patterns [13, 9, 15]. The fundamental mechanisms underlying bioconvection have been extensively modeled through continuum approaches [16, 25, 8], incorporating momentum and mass conservation equations coupled with equations governing the orientation and distribution of swimming cells. Early work by Kessler [12, 24] and later developments by Pedley and Kessler [9] established the theoretical framework for understanding how gyrotaxis and bottom-heaviness drive these large-scale convective phenomena.
Recent studies have demonstrated that cell geometry, fluid depth, and boundary conditions critically influence the formation and stability of bioconvection patterns [18, 26, 27, 28, 29, 30]. While extensive research has explored bioconvection in simple geometries, the effects of more intricate internal structures, such as spiral-shaped boundaries, remain largely unexplored. Spiral geometries are particularly intriguing because they introduce geometrical singularities—regions of sharp curvature and varying local confinement—that can profoundly modulate cell accumulation, flow patterns, and instabilities. Additionally, the presence or absence of an open fluid–air interface significantly impacts bioconvection dynamics, primarily by regulating oxygen exchange between the suspension and the environment. Recent work by Fragkopoulos et al.[31] has shown that metabolic activity—closely linked to both light intensity and oxygen availability—directly modulates microbial motility and fundamentally governs the emergence and stability of coherent flow structures. These findings underscore the pivotal role of metabolic constraints, associated with light and oxygen dynamics, in shaping the spatiotemporal evolution of bioconvection patterns[13, 19, 31, 32].
In this study, we investigate the formation and evolution of bioconvection patterns in suspensions of Chlamydomonas reinhardtii confined within spiral-shaped boundaries, under controlled red-light illumination to minimize phototactic effects. By systematically varying the boundary conditions—from open liquid–air interfaces to fully confined, air-impermeable geometries—we investigate how geometric confinement and oxygen depletion shape the spatiotemporal organization and characteristic wavelengths of bioconvection patterns. Through a combination of detailed experimental approaches—including top-view imaging, space–time analyses, and Fourier transforms—we show that confinement and oxygen depletion induce pronounced transitions in pattern morphology, resulting in the spontaneous reorganization of bioconvection patterns under oxygen-limited conditions. Complementary three-dimensional numerical simulations, based on the incompressible Navier–Stokes equations coupled with active-swimmer models, successfully reproduce the primary stages of pattern formation and hydrodynamic instability. However, these simulations do not capture the oxygen-dependent transitions observed experimentally. Together, our findings demonstrate that variations in confinement geometry and metabolic conditions can significantly alter collective microbial behavior, underscoring a strong interplay between microscale motility, environmental constraints, and emergent fluid flows.
2 Materials and Methods
2.1 Cell culture
[…]
2.2 Experimental setup
To investigate the influence of curved boundaries on bioconvection patterns in CR suspensions, we fabricated molds from acrylic (PMMA)—a material impermeable to air—with different internal geometries, including spiral-shaped and circular designs. Each design was examined under either open (aerobic) or sealed (anaerobic) top boundary conditions to systematically probe the effects of oxygen availability. The experimental setup was carefully leveled and filled with a thin layer of cell suspension (typically 5 mm deep) at a concentration of approximately $10^{7}$ cells/mL. All experiments were conducted inside a dark enclosure to eliminate phototactic responses triggered by ambient light. To further suppress phototactic effects during imaging and pattern development, the suspension was illuminated exclusively from above using a red LED light source centered at 625 nm (LUMIMAX, LSR24-R), with an optical diffuser positioned between the light source and the sample to ensure uniform illumination. Previous studies have demonstrated that the phototactic sensitivity of CR cells declines sharply at longer red wavelengths [35, 36, 37]. Therefore, illumination at red light effectively eliminates phototactic steering, ensuring that the observed bioconvection patterns primarily result from hydrodynamic instabilities and gravitactic behavior.
2.3 Image capture and analysis
For each experiment, over 1000 images were acquired at 1-second intervals using a dark-field imaging setup. In experiments requiring extended observation, patterns were recorded continuously for up to two hours at 1fps. The red LED illumination intensity, measured at the location of the PMMA mold, was maintained at approximately $5\times 10^{17}$ photons m-2 s-1. In the raw dark-field images, regions of high cell density appeared dark, whereas regions of low cell density appeared bright. To quantify the dominant wavelength of the emerging bioconvection patterns, Fourier analysis was performed on space–time plots using custom MATLAB scripts. Prior to analysis, image preprocessing was performed using FIJI. Initially, the images were inverted in intensity, and the mean image was subtracted from the stack to enhance the signal-to-noise ratio. In the resulting 16-bit post-processed images, regions of high cell density appeared bright (visualized in pseudo-green), whereas regions of low density appeared dark.
2.4 Model
Bioconvection, the spontaneous formation of large-scale patterns in suspensions of swimming microorganisms, arises from the interplay between individual cell motility, fluid mechanics, and external fields such as gravity [25, 38, 15, 14, 39, 40, 8]. Most theoretical models of bioconvection aim to link macroscopic flow fields to microscopic swimming behaviors, often under the assumption that direct cell-cell interactions are negligible due to the dilute nature of the suspension. In these systems, the Reynolds number associated with individual swimmers is extremely low (around $10^{-3}$), reflecting viscous-dominated motion at the microscale. However, the collective flows generated by bioconvection can reach much higher Reynolds numbers, often on the order of $10$, highlighting a dramatic amplification of fluid motion through coordinated swimming [28, 8, 14, 4].
Table 1: Parameters used in simulations of bioconvection patterns.
Mean concentration $\bar{c}$                            $4\times 10^{6}$ cells/ml
Fluid depth $L$                            5 mm
Fluid density $\rho_{0}$                            997 kg/m3
Cell density $\rho$                            1050 kg/m3
Specific gravity $\Delta\rho/\rho=\frac{\rho-\rho_{0}}{\rho}$           0.05
Cell radius $a$                            $5\times 10^{-6}$ m
Center of gravity offset $h$                            $10^{-5}$ cm
Volume per cell $\beta$                            $5\times 10^{-16}$ m3
Swimming speed $W_{c}$                         $10^{-4}$ m s-1
Sedimentation speed $W_{s}$                         $6\times 10^{-6}$ m s-1
Diffusivity of cells $D$                           $5\times 10^{-8}$ m2 s-1
Kinematic viscosity $\nu$                            $8.9\times 10^{-7}$ m2 s-1
Gyrotactic reorientation parameter $B$                            1 s
Scaled swimming speed $V_{c}=W_{c}L/D$           10
Schmidt number $S_{c}=\nu/D$               17.85
Gyrotaxis number $G=BD/L^{2}$           $2\times 10^{-3}$
Pseudo-Rayleigh number $R_{a}=\frac{\bar{c}\beta\Delta\rho gL^{3}}{(\rho\nu D)}$           2920
We model the collective dynamics of a dilute suspension of CR cells using a continuum framework based on earlier studies [25, 8, 16]. The model couples the incompressible Navier–Stokes equations with a conservation equation for the cell concentration and an orientation dynamics equation accounting for gyrotaxis:
$\nabla\cdot\bm{U}=0,$ (1)
$\rho_{0}\left(\frac{\partial\bm{U}}{\partial t}+(\bm{U}\cdot\nabla)\bm{U}\right)=-\nabla p-cg\Delta\rho\,\hat{z}+\nabla\cdot(2\mu\bm{E}+\bm{\Sigma}),$ (2)
$\frac{\partial c}{\partial t}=-\nabla\cdot\left[c\left(\bm{U}+W_{c}\bm{p}-W_{s}\hat{z}\right)-D\nabla c\right],$ (3)
$\frac{\partial\bm{p}}{\partial t}=\frac{1}{2B}\left[\hat{z}-(\hat{z}\cdot\bm{p})\bm{p}\right]+\frac{1}{2}\bm{\omega}\times\bm{p},$ (4)
where $\bm{U}$ is the fluid velocity, $\bm{\omega}=\nabla\times\bm{U}$ is the vorticity field, $\mu$ is the fluid viscosity, $\rho_{0}$ is the fluid density, $p$ is the pressure, $c$ is the local cell concentration, and $\bm{p}$ denotes the local swimming direction of the cells. The stress tensor $\bm{E}=(\nabla\bm{U}+\nabla\bm{U}^{T})/2$ represents the rate of strain, and $\bm{\Sigma}$ is the active stress generated by the swimming cells, modeled as $\bm{\Sigma}=cS(\langle\bm{p}\bm{p}\rangle-\bm{I}/3)$ [41, 8, 16, 4]. Here $\bm{I}$ is the unity tensor, $S$ is the stresslet strength and is positive for CR cells which are ’pullers’ and negative for sperm cells which are pushers. The buoyancy term accounts for the density difference $\Delta\rho=\rho-\rho_{0}$ between cells and the surrounding fluid, and the Boussinesq approximation is employed [25, 16, 14]. Given that the characteristic timescale of bioconvection is considerably shorter than that of cellular growth, the conservation of the cell population is described by Eq. 3. This equation incorporates advection by both the fluid velocity and the active swimming motion at speed $W_{c}$, sedimentation under gravity at velocity $W_{s}$, and random motility modeled as an effective diffusion process with diffusivity $D$ [16, 8, 28]. Orientation dynamics (Eq. 4) describes the balance between gravitational torque, which favors upward orientation, and viscous torque arising from fluid shear, resulting in gyrotactic reorientation over a characteristic timescale given by $B=\alpha_{\perp}/2hmg,$ where $\alpha_{\perp}=8\pi\mu a^{3}$ is the rotational drag coefficient for a spherical cell of radius $a$, $h$ is the center-of-gravity offset within the cell, $m$ is the cell mass, and $g$ is the gravitational acceleration [8, 12]. This framework assumes the suspension is sufficiently dilute such that cell-cell hydrodynamic or steric interactions are negligible. Each cell is modeled as a sphere of volume $\beta$, swimming along its axis $\bm{p}$. Table 1 summarizes the parameters used in the numerical simulations, selected based on experimental measurements of CR cells and previous bioconvection studies [15, 4]. Numerical solutions of the coupled system were obtained using COMSOL Multiphysics (version 6.3), enabling us to reproduce key features of the experimentally observed bioconvection patterns.
3 Results
3.1 Spiral-shaped boundaries induce large-scale rotational patterns
Figure 1: Representative top-view images of bioconvection patterns formed by CR cells confined within a spiral-shaped boundary with an open liquid-air interface. (A) Side-view schematic illustrating the experimental setup, showing a fluid depth of 5 mm, approximately 1 mm greater than the spiral boundary height. (B, D) Initial pattern formation at the fluid-air interface, where cells accumulate due to negative gravitaxis, forming lattice-like clusters. Over time, these clusters become increasingly disordered (F, H). Space–time plots (C, G, K) generated by stacking light intensity profiles along trajectories marked in panels B, F and J reveal pronounced rotational dynamics: inward-directed near the spiral core (K) and bidirectional rotation farther from the center (G). Fourier analysis (E, I) confirms a characteristic pattern wavelength of approximately 4–5 mm. See also Video 1 for the full dynamic evolution.
Figure 2: Additional representative images illustrating bioconvection pattern formation under similar experimental conditions to Fig. 1. Initially, cells accumulate at the fluid-air interface, forming distinct spiral-shaped regions of high cell density (B), which subsequently evolve into discrete, lattice-like structures. These structures appear as dense cell clusters (black in raw images, visualized as green spots in the post-processed image, panel C, separated by brighter, interconnected net-like regions of lower cell concentration (E). Space–time plots reveal inward-directed rotational motion near the spiral core (D), whereas areas farther from the center exhibit rotational motion in both clockwise and counterclockwise directions (G). Fourier analysis (I) confirms the characteristic wavelength of the bioconvection patterns to be approximately 5 mm. See also Video 2 for the full dynamic evolution.
Two representative top-view of bioconvection patterns formed by CR cells in a system with a spiral-shaped boundary and an open liquid-air interface is shown in Figs. 1 and  2. In our experiments, the fluid depth is approximately 5 mm, roughly 1 mm greater than the height of the spiral boundary (Figs. 1A). Starting from a homogeneous suspension, cells swim upward due to negative gravitaxis, leading to a lower concentration of cells directly above the spiral boundary compared to the deeper surrounding areas. This results in a distinctive spiral-shaped cell accumulation at the fluid-air interface in areas between the spiral boundary (see Fig. 2B), which is susceptible to bioconvection instability. Consequently, this spiral arrangement fragments into smaller clusters, forming a lattice-like patterns characterized by distinct dark spots, which correspond to regions of high cell density (visualized as green in post-processed images). These dense clusters are separated by brighter, interconnected, net-like structures representing regions of lower cell density (appearing dark in post-processed images). Space-time plots generated along multiple trajectories across different regions of the setup reveal a pronounced large-scale rotational motion in the bioconvection patterns. Notably, near the center of the spiral, we dominantly observe a directed inward rotation toward the spiral core (Fig. 1K and Fig. 2D). In contrast, farther from the center, the rotational behavior becomes more complex, with bioconvection patterns exhibiting both clockwise and counterclockwise motion (Fig. 1G and Fig. 2G). Fourier analysis performed along various trajectories within the setup indicates that the characteristic wavelength of the patterns is approximately 4-5 mm measured in different parts of the channel (Fig. 1E, I and Fig. 2I).
3.2 Oxygen deficiency trigger drastic pattern transitions in confined setups
Figure 3: Evolution of bioconvection patterns under confined conditions with a closed top boundary. (A) Experimental setup with a transparent PMMA wall placed above the spiral structure, maintaining a 1 mm gap. (B–D) Initial lattice-like arrangement of high-density cell clusters (visualized as green spots in post-processed images) separated by lower-density regions. (F–H) Transition phase characterized by fragmentation and the formation of a disorganized, branched cellular network. (J–L) Reorganization into a new, stable pattern with a reduced wavelength. (E, I, M) One-dimensional Fourier transforms of light intensity profiles, extracted along the blue dashed lines in panels D, H, and L, reveal a reduction in the dominant wavelength from approximately 6 mm to 4 mm. See also supplementary Video 3 for the full dynamic evolution.
We modified the boundary conditions by placing a solid transparent PMMA wall above the spiral structure, maintaining a 1 mm gap between the spiral wall and the top boundary. The PMMA-based chamber was completely filled with the cell suspension, ensuring the elimination of any air gaps or bubbles (Fig. 3A). Under these confined conditions, we observed a pronounced transition in the bioconvection patterns over time, which we attribute to oxygen depletion. To characterize this transition, we carried out frame-by-frame analysis, constructed detailed space-time plots, and applied one-dimensional Fourier transforms (1D-FT) to quantify changes in the pattern wavelengths. We measured the wavelengths at three critical stages: before the transition, at its onset, and after the transition. Long-duration recordings (approximately two hours) revealed a significant reduction in the characteristic wavelength, accompanying a dramatic reorganization of the bioconvection patterns.
Similar to the setup with an open liquid-air interface, the bioconvection patterns initially emerged from a homogeneous suspension as localized accumulations of cells arranged in a lattice-like structure. These patterns were characterized by distinct dark regions, corresponding to areas of high cell density (visualized as green spots in post-processed images), separated by brighter regions of lower cell concentration (appearing dark in processed images) (Fig. 3B–C and Video 3). This lattice arrangement persisted for approximately 30–40 minutes before undergoing a drastic transition (Fig. 3F-M). The transition was marked by a temporary weakening or disappearance of the regular pattern, followed by the formation of a disorganized, branched network of cells at random locations (Fig. 3F–I), which rapidly percolated throughout the entire suspension, resulting in a brief intermediate phase of disorganization. Following this phase, the cells spontaneously reorganized into a new, stable, and more regular pattern with a smaller characteristic wavelength (Fig. 3J–M), which remained steady for over an hour as long as the cells retained their motility. Note that the upper portion of the transitional zone—highlighted by a red box in Fig. 3D—exhibits a temporal delay, revealing that the transition event propagates gradually across the suspension.To quantify the associated changes in the pattern wavelength, we applied one-dimensional Fourier transforms to the light intensity profiles extracted from the space-time plots at the selected time points. The results, presented in Fig. 3E, I, and M, reveal a clear reduction in the dominant wavelength from approximately 6 mm to 4 mm during the course of the transition.
Figure 4: Bioconvection pattern evolution under fully confined conditions with no fluid gap. (A) Experimental setup where the top PMMA wall directly contacts the spiral-shaped boundary, eliminating the 1 mm fluid gap present in the previous configuration. (B–D) Sequential top-view images showing the disappearance or significant weakening of the initial bioconvection pattern, followed by reorganization into a new, stable pattern with a smaller characteristic wavelength. (E) Space–time plot illustrating the transition, highlighting two distinct regions with different stripe spacings, separated by a prolonged intermediate phase of disorganization. The longer transition time is attributed to reduced cell-to-cell communication caused by the absence of the fluid gap. Fourier analysis confirms a decrease in the dominant wavelength from approximately 5 mm to 3 mm (panel D); see also Video 4.
Figure 5: Bioconvection pattern evolution under confined conditions with a 1 mm fluid gap (as sketched in Fig. 3A) in a PMMA-made setup containing circular obstacles. (A–C) Sequential top-view images showing the reorganization of bioconvection patterns, with a new pattern featuring a smaller characteristic wavelength emerging after approximately 50 minutes. (D–F) Corresponding processed images of the snapshots shown in panels A–C. (G-I) Space–time plots along the trajectories indicated in panels D–E, revealing the transition, characterized by two distinct regions with different stripe spacings; see also supplementary Video 5.
Building on these observations, we conducted additional experiments under confined conditions, modifying the setup so that the top PMMA wall was in direct contact with the upper edge of the spiral-shaped boundary, thereby eliminating the 1 mm fluid gap present in the previous configuration (Fig. 4A). Sequential top-view images captured at different stages of bioconvection pattern evolution are shown in Fig. 4B–C. Under these fully confined conditions, the intermediate phase—during which the patterns disappeared or substantially weakened—was significantly prolonged, lasting approximately 50 minutes before the system reorganized into a new, stable configuration with a shorter characteristic wavelength (Fig. 4D). This prolonged transition and subsequent reorganization are clearly illustrated in the space–time plot (Fig. 4E), which reveals two distinct regions with different stripe spacings corresponding to pre- and post-transition states, separated by an extended intermediate phase characterized by weak or absent pattern formation. We attribute this delay to impaired fluid-mediated communication between cells due to the absence of the fluid gap, which in the previous configuration facilitated more efficient collective reorganization. Despite the slower dynamics, Fourier analysis confirms a reduction in the dominant pattern wavelength from approximately 5 mm to 3 mm following the transition. To further confirm that oxygen depletion is the underlying cause of this transition, we performed a control experiment in which the entire chamber, including the top boundary, was fabricated from air-permeable PDMS. Under otherwise identical conditions, no pattern transitions were observed (Fig. S1), strongly supporting the conclusion that restricted oxygen exchange in PMMA-based confined geometries is the key factor driving the observed drastic pattern reorganization.
To assess the generality of these findings, we extended our experiments to systems incorporating alternative internal geometries. Specifically, we replaced the spiral boundary with an array of circular obstacles while preserving the same confined conditions used in Fig. 3, including the 1 mm fluid gap. As illustrated in Fig. 5, similar transitions in bioconvection patterns were observed, with the structures reorganizing into a new configuration featuring a reduced characteristic wavelength after roughly 50 minutes. These results demonstrate that the key driver of the transition is the oxygen-depletion itself rather than the specific geometry of the internal boundary. Thus, the striking pattern transitions reported here emerge as robust features of confined bioconvection within air-impermeable boundaries, occurring regardless of whether the internal geometry is spiral, circular, or of another structural configuration.
3.3 Simulations
Figure 6: Top-view snapshots of cell density from 3D simulations of bioconvection patterns above the spiral-shaped boundary at $z=4$ mm. Either a zero-shear-stress condition (representing a free liquid-air interface) or a no-slip condition (representing a rigid top wall) is applied at the upper boundary (at $z=5$ mm). Initially, CR cells swim upward (negative gravitaxis) and accumulate near the surface, forming a distinct spiral-shaped cell density pattern (A, D). This structure becomes dynamically unstable, fragmenting into smaller, high-density clusters that evolve into down-welling plumes (see also Figs. 9-10). Notably, fragmented clusters, initially aligned parallel to the spiral walls (B), progressively reorient perpendicular to the boundaries over time (C, F). A comparison between panels B and E shows that fragmentation of the spiral-shaped accumulation occurs earlier in the system with the zero-shear-stress boundary condition. G) The space–time plot along the spiral trajectory in panel C, constructed from long-term simulation results, reveals the inward movement of plumes toward the spiral center, in good agreement with experimental observations. The color map has been slightly adjusted to enhance visualization; see also Videos 6 and 7.
We investigate three-dimensional bioconvection patterns within a shallow suspension of cells contained in a cylindrical chamber. The simulated geometry closely matches our experimental setup and consists of a rigid bottom surface, rigid cylindrical sidewalls, and a rigid spiral-shaped internal wall immersed in the cell suspension. The fluid layer extends approximately 1 mm above the spiral wall, which has a height of about 4 mm. We apply no-slip boundary conditions at all solid surfaces. At the upper boundary, we impose either a zero-shear-stress boundary condition to represent a free liquid-air interface or a no-slip condition to mimic a rigid top boundary in confined geometry experiments. Additionally, we ensure no flux of cells occurs through the top, bottom, or side boundaries.
Figure 7: Cell density distribution at the bottom of the setup ($z=0$) from 3D simulations of bioconvection patterns. Snapshots at $t=100$ s reveal that the cell concentration at the bottom is approximately half that near the surface (see Fig. 6), and the cell clusters remain more diffuse and unlike the near-surface structures, clusters at the bottom do not exhibit reorientation perpendicular to the spiral boundaries, highlighting the depth-dependent development of bioconvection patterns; see Videos 8 and 9 .
Figure 6 shows a representative top-view of bioconvection patterns that form directly above the spiral-shaped boundary at a height of $z=4$ mm, where $z=0$ corresponds to the bottom of the chamber and $z=5$ mm marks the liquid–air interface. At early times, CR cells exhibit negative gravitaxis, swimming upward against gravity and accumulating near the free liquid-air interface (see also the side view in Fig. 9 and the three-dimensional density profile in Fig. 8A). Owing to the shallower fluid layer directly above the spiral walls—approximately 1 mm compared to 5 mm in surrounding regions—the cell concentration at the surface is initially lower above the spiral boundaries than in adjacent areas. This variation in fluid depth seeds the formation of a distinct spiral-shaped accumulation pattern between the spiral walls, as shown in Fig. 6A, D and Fig. S2A, D, in agreement with experimental observations (Fig. 2B). However, this spiral-shaped accumulation is dynamically unstable due to density inversion, as CR cells are approximately 5% denser than water. As a result, the initially continuous spiral structure fragments into smaller, high-density clusters (Figs. 6B, E), which subsequently evolve into distinct downwelling plume structures (Fig. 9). The formation of these downwelling plumes is accompanied by the emergence of surface flows, as illustrated in Fig. 8B. Notably, the fragmented clusters at $z=4$ mm, which initially align parallel to the spiral walls, progressively reorient perpendicular to the boundaries over time, as depicted in Figs. 6C and F.
Figure 8: (A) Three-dimensional view of the developing bioconvection patterns inside the setup. (B) Zoomed-in top view of the cell density color map and surface flow field with streamlines at $z=4$ mm, highlighting the formation of downwelling plumes. (C) Vertical cut-plane used to extract the density profiles and flow fields shown in Figs. 910; see also Video 10.
Figure 9: Cell density distribution and corresponding flow field with streamlines extracted along a vertical cut plane passing through the center of the spiral-shaped boundary (as defined in Fig. 8C). Panel B shows a zoomed-in view of the region highlighted by the dashed red box in panel A, revealing the formation of large-scale vortices and strong downward plumes, with fluid velocities reaching approximately 1 mm/s—nearly an order of magnitude greater than the typical swimming speed of individual CR cells. The dashed orange box in panel A indicates the region further magnified in Fig. 10; see Video 11.
Figure 10: Cell density distribution and corresponding iso-density contour lines during the formation of downwelling plumes, showing zoomed-in views of the region highlighted by the orange box in Fig. 9 at successive time points.
This reorganization of cell clusters near the surface raises the question of how the underlying bioconvection structures develop throughout the depth of the suspension. To investigate this, we examine the cell density distribution at the bottom of the setup. Figure 7 shows the cell density color map at $z=0$. Interestingly, comparing the snapshots at $t=100$ sec with the corresponding near-surface patterns in Fig. 6 reveals that not only the cell concentration is approximately two times lower at the bottom, but also that the cell clusters remain more diffuse and are not oriented perpendicular to the spiral boundary, in contrast to the reorganization observed near the surface in Figs. 6C and F.
In addition to characterizing the evolving cell density patterns, we investigated the large-scale coherent flow fields generated as downwelling plumes developed within the suspension. To quantify the fluid dynamics associated with bioconvection, velocity fields were extracted along a vertical cut-plane intersecting the center of the spiral-shaped boundary, as defined in Fig. 8C. Figure 9 presents the cell density distribution together with the corresponding velocity field and overlaid streamlines in this plane. To further illustrate the onset of flow instabilities, Fig. 10 displays iso-density contour lines highlighting the emergence and structure of the developing plumes. Consistent with previous observations reported in the literature, we find the formation of vortices featuring strong downward flows concentrated within the plumes, with velocities reaching approximately 1 mm/s—nearly an order of magnitude greater than the typical swimming speed of individual CR cells. These results highlight how collective cell motion dramatically enhances fluid transport far beyond the capabilities of isolated swimmers.
4 Discussion and Conclusions
Our experimental observations reveal striking transitions in bioconvection patterns driven by oxygen depletion, highlighting the critical interplay between boundary conditions, oxygen availability, and microbial collective behavior. Under confined conditions—achieved by placing a transparent PMMA top wall 1 mm above the spiral-shaped boundary—we initially observed stable lattice-like bioconvection patterns, closely resembling those seen in systems with an open liquid-air interface. As the cells gradually consumed the dissolved oxygen, these regular patterns weakened significantly or temporarily disappeared, followed by a spontaneous reorganization into a new, stable arrangement with a markedly smaller wavelength. The transition was initiated at random locations and propagated across the suspension, indicating that localized oxygen depletion progressively modulated cell motility and metabolic activity, thereby driving the onset of collective reorganization. When confinement was further intensified—by eliminating the fluid gap entirely—we observed even more pronounced and prolonged transitions. Under these fully confined conditions, the intermediate phase, during which the bioconvection patterns disappeared, extended to approximately 45 minutes before the system spontaneously reorganized into a new configuration with a reduced wavelength. Reorganization typically began at random locations and gradually percolated throughout the entire suspension. Although oxygen levels are ultimately expected to deplete uniformly throughout the suspension, early-stage spatial variations in cell concentration leads to heterogeneous local depletion rates, causing pattern reorganization to initiate at random locations across the system. Collectively, these results demonstrate that subtle changes in confinement geometry and oxygen availability can profoundly influence the dynamics, timescales, and structural evolution of bioconvection patterns, offering promising strategies for manipulating microbial collective behavior through environmental and geometric control.
Our findings are consistent with the recent study by Fragkopoulos et al. [31], who demonstrated that under anaerobic conditions, confined suspensions of photosynthetically active microorganisms exhibit reversibly switchable coherent flows. In their experiments, lowering the red-light intensity under anaerobic conditions—thereby reducing photosynthetic activity—led to a marked decrease in cell swimming speeds (from approximately 100,$\mu$m,s-1 to 40,$\mu$m,s-1) and initiated the emergence of spatially regular, temporally stationary bioconvective plumes within an otherwise stable suspension exhibiting an inverted sedimentation profile. While Fragkopoulos et al. systematically varied the light intensity to modulate metabolic activity, in our experiments the red-light intensity was held constant at approximately $5\times 10^{17}$ photons  m-2 s-1 throughout. Instead, oxygen availability was modulated purely through physical confinement. We propose that this confinement-induced oxygen depletion similarly drives a metabolic transition from aerobic to anaerobic conditions, ultimately leading to the bioconvection pattern transitions observed in our system. Thus, both work independently support the conclusion that metabolic activity and oxygen dynamics are key regulators of collective behavior in motile microbial suspensions and further emphasizes that geometric confinement alone, by limiting oxygen replenishment, can drive dramatic reorganizations of bioconvection patterns even in otherwise constant environmental conditions.
Furthermore, an earlier study by Kage et al. [42] reported drastic reorganization of bioconvection patterns in confined CR suspensions under red-light illumination, which they primarily attributed to gyrotactic reorientation—namely, changes in the balance between gravitational and viscous torques experienced by the cells. In that study, the authors also observed that pattern wavelength shifts correlated with illumination intensity, noting that red light predominantly influenced the steady-state patterns after the transition, rather than affecting the initial onset or the pre-transition steady states. They proposed that these changes could be linked to alterations in cellular photosynthetic activity. In connection with these findings, our results—obtained under constant red-light illumination—highlight that confinement, and the associated oxygen depletion, likely inducing a metabolic shift from aerobic respiration to photosynthesis, serves as a critical driver of the observed pattern transitions. Notably, in our experiments, no pattern transitions were observed under open liquid–air interface conditions, nor in control experiments where the entire setup—including the top boundary—was fabricated from oxygen-permeable PDMS (Fig. S1). In contrast, striking pattern reorganizations consistently emerged under confined PMMA-based configurations with restricted oxygen replenishment, underscoring the critical role of oxygen availability in driving the observed transitions.
To better understand the mechanisms underlying the pattern transitions observed in our experiments, we performed numerical simulations based on standard continuum models of bioconvection. These models, which couple the incompressible Navier–Stokes equations with cell conservation and gyrotactic reorientation [38, 43], have been widely used to describe bioconvection instabilities. In these frameworks, the onset of bioconvection is predicted when the Rayleigh number exceeds a critical threshold, with characteristic pattern wavelengths determined by the fastest-growing unstable modes. However, it is important to recognize the inherent limitations of these classical models. They typically assume constant swimming speeds and negligible cell–cell interactions under the dilute-suspension approximation. As noted previously [43, 15], in the nonlinear stages of bioconvection, local cell concentrations within dense, downwelling plumes can significantly exceed the initial mean value, challenging the validity of the dilute-suspension assumption and potentially altering hydrodynamic and collective behaviors. In our confined system, the observed pattern transitions—characterized by the temporary disappearance and subsequent reorganization into smaller-wavelength structures—cannot be explained solely by these classical bioconvection models. Even when our numerical simulations incorporated no-slip boundary conditions to emulate experimental confinement, the standard framework was only able to reproduce the initial formation and fragmentation of spiral-shaped cell accumulations into downwelling plumes; it failed to capture the striking reorganizations into new, stable patterns observed experimentally. This discrepancy highlights the need to extend standard bioconvection theories to account for additional factors, notably oxygen dynamics. Our experiments strongly suggest that oxygen depletion—and the associated transition from aerobic to anaerobic metabolism, which was shown to modulate microbial motility in Ref. [31]—plays a critical role in driving these transitions. Indeed, similar to the modeling of oxytactic bacterial suspensions [15, 44, 45, 46, 47], it may be necessary to incorporate an explicit oxygen concentration field $C_{oxy}(\bm{x},t)$, coupled to microbial motion. In these models, the average cell swimming velocity arises as a result of chemotaxis and is assumed to be directly proportional to the gradient of the oxygen concentration:
$W_{c}=\chi\nabla C_{oxy},$ (5)
where $\chi$ is the chemotactic sensitivity coefficient. Consequently, a further evolution equation for $C_{\text{oxy}}$ must be introduced, incorporating diffusion, advection by the fluid, and consumption by the microorganisms:
$\frac{\partial C_{\text{oxy}}}{\partial t}+\mathbf{U}.\nabla C_{\text{oxy}}=D_{\text{oxy}}\nabla^{2}C_{\text{oxy}}-\gamma c,$ (6)
where $D_{\text{oxy}}$ denotes the molecular diffusivity of oxygen (typically much larger than the effective diffusivity of the cells), $\gamma$ is the oxygen consumption rate per cell, and $c$ represents the local cell concentration. To fully capture the complex spatiotemporal dynamics and pattern transitions observed in confined active suspensions, future bioconvection models must explicitly incorporate oxygen transport and depletion. Moreover, the simple linear chemotactic response assumed in Eq. (5) may require refinement, as cellular swimming behavior could vary nonlinearly with oxygen concentration, particularly under conditions of severe depletion or strong spatial heterogeneity. Incorporating metabolically regulated motility responses will thus be critical for accurately describing the rich transitions uncovered in our experiments. Finally, potential crowding effects at elevated cell densities should also be considered to fully account for the intricate interplay between metabolic activity, confinement, and collective flow phenomena observed in our system.
In summary, our study highlights the pivotal role of metabolic activity and oxygen dynamics in governing collective microbial behavior and demonstrates how environmental design—through boundary geometry and oxygen control—can be leveraged to manipulate microbial self-organization and fluid transport in active matter systems. More broadly, these findings offer new strategies for directing microbial collective flows and structuring living fluids through targeted environmental and geometric interventions, with potential applications in bioreactors, synthetic ecosystems, and the design of bioinspired active materials.
Acknowledgements
We sincerely thank Professors L. Dede, G. Shubeita, A. Narayanan, M. Ali, A. Pumir, H. de Maleprade, F. Picella, E. Bodenschatz and S. Ramaswamy for their invaluable insights and stimulating discussions. We are also grateful to Dr. Malavath for his dedicated support. Special thanks are extended to the Core Technology Platform (CTP) at NYU Abu Dhabi for their exceptional assistance, particularly to V. Dhanvi and J. Govindan for their outstanding work in fabricating the acrylic molds. Finally, we acknowledge the High Performance Computing (HPC) facility at NYU Abu Dhabi for providing critical computational resources and data storage support essential to this study.
Author contributions
S.G., S.O.A. and T.D. conducted the experiments, I.G., S.V.R.A., A.B. and A.G. designed and carried out the numerical simulations. A.G. conceived the idea, designed the experiments, analyzed the data, and wrote the manuscript.
Data and materials availability All data needed to evaluate the conclusions in the paper are present in the paper and/or the Supplementary Materials. Additional data related to this paper may be requested from the authors.